A body of mass 5 Kg is raised vertically to a height of 10 m by a force of 170 N. The final velocity of the body is
A. 15 m/s B. 17 m/s C. 20 m/s D. 22 m/s

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
F=MA
170 N = 5 kg . a
a = 34 m/s^2
But this a has to also overcome 10 m/s^2 of gravity
Therefore a = 24 m/s^2
Now v^2-u^2=2as
v^2 = 2 a s = 2* 24*10 = 480
Therefore v = sqrt (480) m/s=22m/s
Thanks and Regards
Shaik Aasif
askIITians faculty